iosangel:

Graph the equation on the coordinate plane y=1/3x

iosangel:

Vocaloid:

well an easy way to start is by plugging in some x values x = 0 gives us y = (1/3)*0 = 0, right? so (0,0), or the origin, is a point

Vocaloid:

x = 1 gives us (1/3)(1) = 1/3, right? so (1,1/3) would be the next point. you can keep going up with x values like 2, 3, 4... etc. or go down to -1, -2, -3, etc.

iosangel:

Like this?

Vocaloid:

1/3 is a bit lower than that haha if it helps, try graphing x = 3 (which gives you 3,1) and then connecting (3,1) and (0,0)

Vocaloid:

if you only graph every three x-values (so x = -3, x = 0, x = 3, x = 6, etc.) you only have to deal with whole numbers